OpenStudy (anonymous):

A machine needs 2 of its gears replaced. You find 2 gears, one manufactured 5 years ago and the other manufactured 10 years ago. The specifications sheet says that the chance of failure is 5% after 5 years and increases to 8% after 10 years. If the gears on the machine are in no way linked to each other, what are the chances that both will fail?

OpenStudy (epoweritheta):

chance of failure = 5/100 *8/100=4/1000=0.004